OpenStudy (destinykiara99):

Simplify the expression. cos^2x + sin^2x / cot^2x-csc^2x

OpenStudy (destinykiara99):

csc x -1 1 sec x

jhonyy9 (jhonyy9):

hope you know that cos^2 x +sin^2 x = 1 yes ?

jhonyy9 (jhonyy9):

csc x = ? do oyu know it ?

OpenStudy (destinykiara99):

No, I have no idea.

jhonyy9 (jhonyy9):

csc x = 1/sin x

OpenStudy (sshayer):

\[\cos ^2x+\sin ^2x=1\] \[\csc ^2x-\cot ^2x=1\]